OpenStudy (anonymous):

x+5=7 Ok what could I do Farter to Isolate the variable? 1. Add 5 to both sides 2.Subtract 7 from both sides 3.Subtract 5 from both sides

OpenStudy (anonymous):

I Know it cant be 2. Because then it would turn the 5 in to a negative so it must be 3 Right?

OpenStudy (luffingsails):

If I wanted to "undo" the adding of the five on the left... I would subtract the 5 from each side. Yes, 3 is right.

OpenStudy (anonymous):

x+5=7 -5 -5 -------- x=2 Right?

OpenStudy (luffingsails):

Yes. Good job.

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

3 is the correct option